# LatentExplainer: Explaining Latent Representations in Deep Generative Models with Multi-modal Foundation Models

### v3

@@ Abstract
Deep generative models like VAEs and diffusion models have advanced various generation tasks by leveraging latent variables to learn data distributions and generate high-quality samples. Despite the field of explainable AI making strides in interpreting machine learning models, understanding latent variables in generative models remains challenging. This paper introduces LatentExplainer, a framework for automatically generating semantically meaningful explanations of latent variables in deep generative models. LatentExplainer tackles three main challenges: inferring the meaning of latent variables, aligning explanations with inductive biases, and handling varying degrees of explainability. By perturbing latent variables and interpreting changes in generated data, the framework provides a systematic approach to understanding and controlling the data generation process, enhancing the transparency and interpretability of deep generative models. We evaluate our proposed method on several real-world and synthetic datasets, and the results demonstrate superior performance in generating high-quality explanations of latent variables.
@@ 1 Introduction
Deep generative models, such as Variational Autoencoders (VAEs) Kingma and Welling (2013) and diffusion models Rombach et al. (2022), have become a state-of-the-art approach in various generation tasks Razavi et al. (2019); Ho et al. (2022); Yang et al. (2023). This method effectively leverages latent variables to learn underlying data distributions and generate high-quality samples by capturing the underlying structure of high-dimensional data in a low-dimensional semantic space Vahdat et al. (2021). To enhance performance, inductive biases are often enforced over latent variables. For instance, disentanglement is a rule of thumb which enforces orthogonality among different latent variables Ding et al. (2020). Moreover, sometimes latent variables can be grouped, leading to combination bias Klys et al. (2018); Guo et al. (2020). More recently, the controllability of deep generative models is desired, where latent variables are bound with specific properties of interest Wang et al. (2024). As the latent variables represent all the information in a lower dimension, they can be considered as an effective abstraction of data into key factors. Hence, it is highly beneficial to interpret the meaning of the latent variables to understand the key factors of the data better and to control the data generation process more effectively. For examples, as shown in Figure 1, by interpreting the latent variables of deep generative models such as VAEs for images, we can elicit key factors of the images for better understanding and control of the data generation.
Figure 1: Latent variables of deep generative models represent data distribution in a low-dimensional space. This figure demonstrates how the interpretation of latent variables’ semantic meaning helps to generate data samples with controllable properties.
@@
Despite the progress in XAI, interpreting latent variables in deep generative models presents significant challenges. First, these variables are not grounded in real-world concepts, and the black-box nature of the models prevents us from inferring the meaning of latent variables from observations. Second, explanations must adhere to the inductive biases imposed on the latent variables, which is essential yet difficult to ensure. For example, in disentangled latent variables, the semantic meanings should be orthogonal. Third, different latent variables have varying degrees of explainability. Some may be trivial to data generation and intrinsically lack semantic meaning. It is crucial to identify which latent variables are explainable and which do not need explanation.
To address the aforementioned challenges, we propose LatentExplainer, a novel and generic framework that automatically generates semantically-meaningful explanations of latent variables in deep generative models. To explain these variables and work around the black-box nature of the models (Challenge 1), we propose to perturbe each latent variable and explain the resulting changes in the generated data. Specifically, we perturb and decode each manipulated latent variable to produce the corresponding sequence of generated data samples. The trend in the sequence is leveraged to reflect the semantics of the latent variable to be explained. To align explanations with the intrinsic nature of the deep generative models (Challenge 2), we design a generic framework that formulates inductive biases on the Bayesian network of latent variable models into textual prompts. These prompts are understandable to humans and large foundation models, hence converting user-provided inductive bias formulas into corresponding prompts that a multimodal large language model can easily understand. To handle the varying degree of explainability in latent variables (Challenge 3), we propose to probe the confidence of the explanations to avoid hallucitations by estimating their uncertainty. This approach assesses whether the latent representations are interpretable and selects the most consistent explanations, ensuring accurate and meaningful interpretation of the latent variables.
@@ 2 Related Work
Image Manipulation. Image manipulation using latent variables in generative models like VAEs and diffusion models is an important technique for modifying and enhancing generated images. A key method is latent traverse, which involves traversing different values of latent variables to achieve diverse manipulations in the generated outputs. This technique allows for precise control over the attributes in generated images, enabling adjustments Chen et al. (2016); Higgins et al. (2017). For example, latent traverse has been effectively employed to disentangle and control various attributes in generated images Brock et al. (2016); Zhu et al. (2016). Although latent traverse is used for visualization and editing purposes, it has not yet been widely explored as a tool for explaining the underlying generative processes in natural language.
Multi-modal Large Language Models. Multi-modal Large Language Models (MLLMs) integrate diverse data modalities, enhancing their ability to understand and generate complex information Yin et al. (2023); Bai et al. (2024). Notable models include CLIP, which learns visual and textual representations Radford et al. (2021), GPT-4o, which extends GPT-4v with better visual capabilities OpenAI (2024), and Gemini that is a family of highly capable multimodal models Team et al. (2023) to perform complex tasks that require understanding context across different modalities.
@@ 3 Preliminaries and Problem Formulation
Inductive Bias in Latent Variables. Inductive biases are commonly imposed on latent variables to enhance the performance and interpretability of deep generative models. These biases can be categorized into several types: Disentanglement Bias: It enforces orthogonality among different latent variables, ensuring that each latent variable captures a distinct factor of variation in the data Ding et al. (2020). Combination bias: Sometimes latent variables are grouped, leading to biases in how they interact and combine to represent complex data structures Klys et al. (2018). Conditional Bias: It emphasizes the relationship between specific properties of interest and the corresponding latent variables Guo et al. (2020). Other Biases: Various other biases can be introduced depending on the specific requirements of the model, such as sparsity or hierarchical organization, to guide the learning process more effectively.
Interpreting latent variables in deep generative models is extremely challenging due to: 1). Their lack of grounding in real-world concepts and the models’ black-box nature, which hinders meaning inference. 2). The necessity for explanations to conform to inductive biases, such as orthogonality in disentangled variables, which is difficult to ensure. 3). The varying degrees of explainability among latent variables, with some being trivial to data generation and lacking semantic meaning, making it essential to identify which variables require explanation.
@@ 4.2.1 Generic Framework
Framework: Our framework proposes a principled, automatic way that translate the mathematical expression to textual prompts. The prompts include adaptive prompts and fixed prompts. The adaptive prompts are converted from the inductive bias formulas. The formula contains mathematical symbols that consist of mathematical variables and mathematical operators. We use the same color to represent the correspondence between mathematical symbols in the formulas and the text in the prompts. The translation mechanism is shown in Table 1.
@@ 4.3 Automatic In-context Prompt Generation
By leveraging these three common inductive biases identified in generative models, we can automatically generate prompts $\mathcal{P}$ that align with $\mathcal{F}$ using in-context learning. The pseudo-code of this section can be found at Algorithm 1.
@@
Finally, the algorithm generates the prompt $\mathcal{P}$ using the formula $\mathcal{F}$, the few-shot examples $\mathcal{H}$, and the gathered semantics (line 8). This step-by-step reasoning process ensures that the generated prompts are contextually relevant and aligned with the underlying formulas, which could reduce hallucination and enhance model performance and interoperability.
Model Dataset BLEU ROUGE-L METEOR SPICE Accuracy
DDPM w/ Inductive Bias Prompts CelebA-HQ 18.5 35.1 32.8 18.8 71.1
LSUN-Church 30.9 44.3 36.4 29.2 78.9
AFHQ-Dog 25.9 38.0 41.5 29.9 78.9
DDPM w/o Inductive Bias Prompts CelebA-HQ 0.2 4.1 4.3 1.8 13.8
LSUN-Church 8.9 28.1 20.6 14.5 55.6
AFHQ-Dog 2.9 13.3 13.8 9.0 39.1
$\beta$-TCVAE w/ Inductive Bias Prompts CelebA-HQ 29.9 55.4 44.3 30.2 83.3
3DShapes 25.4 49.1 32.7 22.8 88.9
dSprites 12.5 37.3 29.9 21.7 73.3
$\beta$-TCVAE w/o Inductive Bias Prompts CelebA-HQ 6.1 32.7 25.3 14.0 50.0
3DShapes 5.4 31.2 22.3 10.1 88.9
dSprites 0.0 18.7 14.4 7.5 13.3
@@
Model Dataset BLEU ROUGE-L METEOR SPICE Accuracy
CSVAE w/ Inductive Bias Prompts CelebA-HQ 25.3 45.3 39.8 27.0 93.3
3DShapes 36.2 43.3 36.6 36.8 80.0
dSprites 16.0 35.9 29.2 21.8 66.7
CSVAE w/o Inductive Bias Prompts CelebA-HQ 12.0 37.1 28.5 17.5 80.0
3DShapes 14.2 28.6 25.1 11.8 50.0
dSprites 0.0 16.4 10.1 7.3 25.0
@@
Model Dataset BLEU ROUGE-L METEOR SPICE Accuracy
Stable Diffusion w/ Inductive Bias Prompts CelebA-HQ 18.5 40.9 33.9 23.9 82.7
LSUN-Church 18.3 38.1 32.5 25.0 97.3
AFHQ-Cat 13.3 29.7 23.6 20.1 88.0
Stable Diffusion w/o Inductive Bias Prompts CelebA-HQ 5.9 24.3 21.1 12.2 45.3
LSUN-Church 10.5 29.8 26.4 17.0 76.0
AFHQ-Cat 10.2 28.6 24.6 17.3 68.0
CSVAE w/ Inductive Bias Prompts CelebA-HQ 24.2 42.9 36.2 19.6 100.0
3DShapes 25.7 39.5 31.6 20.7 73.3
dSprites 21.2 35.0 28.7 20.7 53.3
CSVAE w/o Inductive Bias Prompts CelebA-HQ 6.1 32.7 25.3 14.0 53.3
3DShapes 5.4 31.2 22.3 10.1 80.0
dSprites 0.0 18.7 14.4 7.5 13.3
@@
4.4 Uncertainty-aware Explanations
Uncertainty-aware methods can be applied to large language model responses Lin et al. (2023) and image explanations Zhao et al. (2024). To measure the uncertainty of the responses from GPT-4o, we sampled $n$ times from the GPT-4o to generate the responses =$\mathcal{R}=\{r_{1},r_{2},r_{3},...,r_{n}\}.$ The certainty score of the explanation is the average pairwise cosine similarity of the responses $\mathcal{R}$: $1/C\cdot{\sum}_{i=1}^{n}{\sum}_{j=1,i\neq j}^{n}\text{sim}(r_{i},r_{j})$ where $C=n\cdot(n-1)$. We further use a threshold of the uncertainty score to assess the interpretability of the latent variables. The implementation details are in Appendix B.
5 Experiment
The experiments are conducted on a 64-bit machine with 24-core Intel 13th Gen Core i9-13900K @ 5.80GHz, 32GB memory and NVIDIA GeForce RTX 4090. We use GPT-4o as our MLLM and set $temperature=1$ and $top\_p=1.$ The code is available at  https://github.com/mengdanzhu/LatentExplainer.
5.1 Dataset
We utilized five datasets to evaluate the performance of different generative models under three inductive biases: CelebA-HQ (Karras et al., 2018), AFHQ-Dog (Choi et al., 2020), LSUN-Church (Yu et al., 2015) for the unconditional and conditional diffusion models, and CelebA-HQ (Karras et al., 2018), 3DShapes (Burgess and Kim, 2018), and dSprites (Matthey et al., 2017) for the VAE models. The CelebA-HQ dataset is a high-quality version of the CelebA dataset, consisting of 30K images of celebrities, divided into 28K for training and 2K for testing. The LSUN-Church dataset contains large-scale images of church buildings. The AFHQ dataset includes high-quality images of animals, divided into three categories: cats, dogs, and wild animals. 3DShapes is a synthetic dataset contains images of 3D shapes with six factors of variation: floor hue, wall hue, object hue, object scale, object shape, and wall orientation, divided into 384K for training and 96K for testing. dSprites consists of 2D shapes (hearts, squares, ellipses) generated with five factors of variation: shape, scale, orientation, position X, and position Y, divided into 516K for training and 221K for testing.

### v4

@@ Abstract
Deep generative models like VAEs and diffusion models have advanced various generation tasks by leveraging latent variables to learn data distributions and generate high-quality samples. Despite the field of explainable AI making strides in interpreting machine learning models, understanding latent variables in generative models remains challenging. This paper introduces LatentExplainer, a framework for automatically generating semantically meaningful explanations of latent variables in deep generative models. LatentExplainer tackles three main challenges: inferring the meaning of latent variables, aligning explanations with inductive biases, and handling varying degrees of explainability. Our approach perturbs latent variables, interpreting changes in generated data, and uses multi-modal large language models (MLLMs) to produce human-understandable explanations. We evaluate our proposed method on several real-world and synthetic datasets, and the results demonstrate superior performance in generating high-quality explanations for latent variables. The results highlight the effectiveness of incorporating inductive biases and uncertainty quantification, significantly enhancing model interpretability.
@@ 1 Introduction
Deep generative models, such as Variational Autoencoders (VAEs) Kingma and Welling (2013) and diffusion models Rombach et al. (2022), have become a state-of-the-art approach in various generation tasks Razavi et al. (2019); Ho et al. (2022); Yang et al. (2023). This method effectively leverages latent variables to learn underlying data distributions and generate high-quality samples by capturing the underlying structure of high-dimensional data in a low-dimensional semantic space Vahdat et al. (2021). To enhance performance, inductive biases are often enforced over latent variables. For instance, disentanglement is a rule of thumb which enforces orthogonality among different latent variables Ding et al. (2020). Moreover, sometimes latent variables can be grouped, leading to combination bias Klys et al. (2018); Guo et al. (2020). More recently, the desire for controllability in deep generative models, where latent variables are associated with specific properties of interest, Wang et al. (2024), has given rise to conditional bias. As the latent variables represent all the information in a lower dimension, they can be considered as an effective abstraction of key factors in the data. Therefore, it is critical to develop methods for automatically decomposing and explaining meaningful latent dimension semantics given a pretrained generative model and its inherent inductive biases, as illustrated in Figure 1. Such capabilities are essential for applications in controllable data generation, interpretable model explanations, and natural language understanding.
Figure 1: Illustration of how a pretrained generative model, guided by inductive bias formulas, automatically decodes and interprets meaningful latent dimension semantics. Zoom in for details.
@@
Despite the progress in XAI, interpreting latent variables in deep generative models presents significant challenges. First, these variables are not grounded in real-world concepts, and the black-box nature of the models prevents us from inferring the meaning of latent variables from observations. Second, explanations must adhere to the inductive biases imposed on the latent variables, which is essential yet difficult to ensure. For example, in disentangled latent variables, the semantic meanings should be orthogonal. Third, different latent variables have varying degrees of explainability. Some may be trivial to data generation and intrinsically lack semantic meaning. It is crucial to identify which latent variables are explainable and which do not need explanations.
To address the aforementioned challenges, we propose LatentExplainer, a novel and generic framework that automatically generates semantically-meaningful explanations of latent variables in deep generative models. Specifically, to explain these variables and work around the black-box nature of the models (Solution to Challenge 1), we propose to perturb each latent variable and explain the resulting changes in the generated data. Specifically, we perturb and decode each manipulated latent variable to produce the corresponding sequence of generated data samples. The trend in the sequence is leveraged to reflect the semantics of the latent variable to be explained. To align explanations with the intrinsic nature of the deep generative models (Solution to Challenge 2), we design a generic framework that formulates inductive biases on the Bayesian network of latent variable models into textual prompts. These prompts are understandable to large foundation models and humans. To handle the varying degree of explainability in latent variables (Solution to Challenge 3), we propose to measure the confidence of the explanations by estimating their uncertainty. This approach assesses whether the latent representations are interpretable and selects the most consistent explanations, ensuring accurate and meaningful interpretation of the latent variables.
@@ 2 Related Work
Latent Variable Manipulation and Explanations. Manipulating latent variables in generative models like VAEs and diffusion models is an important technique for editing and enhancing generated images. A key method is latent traverse, which involves traversing different values of latent variables to achieve diverse manipulations in the generated outputs. This technique allows for precise control over the attributes in generated images, enabling adjustments Chen et al. (2016); Higgins et al. (2017). For example, latent traverse has been effectively employed to disentangle and control various attributes in generated images Brock et al. (2016); Zhu et al. (2016). However, latent traverse is often used for visualization and editing purposes. It has not yet been widely explored as a tool for explaining the underlying latent space. Their explanations primarily rely on using predefined training attributes as text labels or manually adding explanations Shen et al. (2020); Esser et al. (2020). Some concept-based models control latent variables in generative models using category concepts to generate data Tran et al. (2022); Bouchacourt et al. (2018). However, these approaches cannot automatically generate free-form textual explanations.
More recently, Multi-modal Large Language Models (MLLMs) integrate diverse data modalities, enhancing their ability to understand and generate complex information Yin et al. (2023); Bai et al. (2024). Notable models include GPT-4o, which extends GPT-4v with better visual capabilities OpenAI (2024), and Gemini that is a family of highly capable multimodal models Team et al. (2023). We envision leveraging MLLMs to automatically generate explanations for latent variables. Our work focuses on: (1) how to develop prompts that are aligned with the underlying generative model and can be easily understood by MLLMs, as well as (2) how to evaluate the quality of the generated explanations for latent variables.
@@ 3 Preliminaries and Problem Formulation
Inductive Bias in Latent Variables. Inductive biases are commonly imposed on latent variables to enhance the performance and interpretability of deep generative models. These biases can be categorized into several types: Disentanglement Bias: It enforces orthogonality among different latent variables, ensuring that each latent variable captures a distinct factor of variation in the data Ding et al. (2020). Combination bias: Sometimes latent variables are grouped, leading to biases in how they interact and combine to represent complex data structures Klys et al. (2018). Conditional Bias: It emphasizes the relationship between specific properties of interest and the corresponding latent variables Guo et al. (2020). Other Biases: Various other biases can be introduced depending on the specific requirements of the model to guide the learning process more effectively.
@@ 4.2.1 Generic Framework
Framework: Our framework proposes a principled, automatic way that translate the mathematical expression to textual prompts. The prompts include adaptive prompts and a fixed ending. The adaptive prompts are converted from the inductive bias formulas. The formulas contain mathematical symbols that consist of mathematical variables and mathematical operators. We use the same color to represent the correspondence between mathematical symbols in the formulas and the text in the prompts. The translation mechanism of adaptive prompts is shown in Table 1.
@@ 4.3 Automatic In-context Prompt Generation
By leveraging these three common inductive biases identified in generative models, we can automatically generate prompts $\mathcal{P}$ that align with $\mathcal{F}$ using in-context learning. The pseudo-code of this section can be found at Algorithm 1 and a more detailed example is at Appendix  C.
@@
Finally, the algorithm generates the prompt $\mathcal{P}$ using the formula $\mathcal{F}$, the few-shot examples $\mathcal{H}$, and the gathered semantics (line 8). This step-by-step reasoning process ensures that the generated prompts are contextually relevant and aligned with the underlying formulas, which could reduce hallucination and enhance model performance.
4.4 Uncertainty-aware Explanations
Uncertainty-aware methods can be applied to large language model responses Lin et al. (2023) and image explanations Zhao et al. (2024). To measure the uncertainty of the responses from GPT-4o, we sampled $n$ times from the GPT-4o to generate the responses =$\mathcal{R}=\{r_{1},r_{2},r_{3},...,r_{n}\}.$ The certainty score of the explanation is the average pairwise cosine similarity of the responses $\mathcal{R}$: $1/C\cdot{\sum}_{i=1}^{n}{\sum}_{j=1,i\neq j}^{n}\text{sim}(r_{i},r_{j})$ where $C=n\cdot(n-1)$. The implementation details is in Appendix B.
5 Experiment
The experiments are conducted on a 64-bit machine with 24-core Intel 13th Gen Core i9-13900K @ 5.80GHz, 32GB memory and NVIDIA GeForce RTX 4090. We use GPT-4o as our MLLM and set $temperature=1$ and $top\_p=1.$ The code is available at  https://anonymous.4open.science/r/LatentExplainer-89AF.
5.1 Dataset
We utilized five datasets to evaluate the performance of different generative models under three inductive biases: CelebA-HQ (Karras et al., 2018), AFHQ (Choi et al., 2020), LSUN-Church (Yu et al., 2015) for the unconditional and conditional diffusion models, and CelebA-HQ (Karras et al., 2018), 3DShapes (Burgess and Kim, 2018), and dSprites (Matthey et al., 2017) for the VAE models. The CelebA-HQ dataset is a high-quality version of the CelebA dataset, consisting of 30K images of celebrities, divided into 28K for training and 2K for testing. The LSUN-Church dataset contains large-scale images of church buildings. The AFHQ dataset includes high-quality images of animals, divided into three categories: cats, dogs, and wild animals. 3DShapes is a synthetic dataset contains images of 3D shapes with six factors of variation: floor hue, wall hue, object hue, object scale, object shape, and wall orientation, divided into 384K for training and 96K for testing. dSprites consists of 2D shapes (hearts, squares, ellipses) generated with five factors of variation: shape, scale, orientation, position X, and position Y, divided into 516K for training and 221K for testing.
Model Method BLEU@4 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$ SPICE $\uparrow$ BERTScore $\uparrow$ BARTScore $\uparrow$
DDPM GPT-4o Baseline 2.90 13.81 11.84 7.08 87.33 -3.32
+ LatentExplainer w/o IB 3.70 14.22 12.17 7.90 87.30 -3.31
+ LatentExplainer w/o UQ 16.81 31.99 29.04 19.79 89.88 -3.08
+ LatentExplainer 25.11 39.11 36.89 25.97 91.19 -2.98
$\beta$-TCVAE GPT-4o Baseline 5.53 27.70 19.43 10.73 88.97 -3.14
+ LatentExplainer w/o IB 3.85 27.71 20.68 10.55 88.93 -3.15
+ LatentExplainer w/o UQ 18.70 41.13 33.84 22.63 91.03 -2.86
+ LatentExplainer 22.63 47.35 35.63 24.88 91.91 -2.80
@@ 5.1 Dataset
Model Method BLEU@4 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$ SPICE $\uparrow$ BERTScore $\uparrow$ BARTScore $\uparrow$
CSVAE GPT-4o Baseline 5.79 28.03 20.24 10.58 89.05 -3.02
+ LatentExplainer w/o IB 8.73 27.32 21.25 12.18 89.16 -3.03
+ LatentExplainer w/o UQ 16.21 35.60 27.58 22.81 90.57 -2.80
+ LatentExplainer 25.85 41.54 35.22 28.51 91.29 -2.74
@@
Model Method BLEU@4 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$ SPICE $\uparrow$ BERTScore $\uparrow$ BARTScore $\uparrow$
Stable Diffusion GPT-4o Baseline 7.07 25.65 21.79 14.72 89.70 -3.27
+ LatentExplainer w/o IB 8.35 26.32 22.94 14.77 89.61 -3.25
+ LatentExplainer w/o UQ 14.85 32.83 27.03 22.02 90.61 -3.12
+ LatentExplainer 16.70 36.32 29.97 23.00 90.97 -3.08
CSVAE GPT-4o Baseline 6.48 22.98 17.03 8.55 88.49 -3.11
+ LatentExplainer w/o IB 9.21 26.68 21.09 11.04 88.98 -3.05
+ LatentExplainer w/o UQ 11.82 29.56 22.07 14.20 88.89 -3.01
+ LatentExplainer 23.70 39.29 32.16 20.33 90.53 -2.89
@@
5.2 Generative Models under Inductive Biases
5.3 Human Annotaions
The ground-truth annotations of explanations are performed by four annotators from the United States and China. All annotators are students that had at least an undergraduate degree. Annotators were presented with the same images and the prompts as MLLM and were asked to annotate the pattern of the images. If there is no clear pattern, just write "No clear explanation". The annotations are then aggregated as references to calculate the automated evaluation metrics.
5.4 Quantitative Analysis
For the quantitative explanation evaluation, we use BLEU Papineni et al. (2002), ROUGE-L Lin (2004), METEOR Banerjee and Lavie (2005), SPICE Anderson et al. (2016), BERTScoreZhang et al. (2019), and BARTScore Yuan et al. (2021) as the automated metrics to assess the generated explanations. BLEU, ROUGE-L, METEOR are n-gram-based metrics that measure the overlap between generated and reference texts. SPICE compares scene graphs derived from the generated and reference texts. BERTScore and BARTScore utilize pre-trained transformer-based language models to compute contextual embeddings of the generated and reference texts. These metrics together provide a comprehensive assessment of both the lexical and semantic quality of the explanations.
To understand the impact of various components in our proposed LatentExplainer, we conducted a comprehensive quantitative analysis across different models (DDPM, $\beta$-TCVAE, Stable Diffusion, and CSVAE) and inductive biases (disentanglement, combination, and conditional). Specifically, we compare the removal of inductive bias prompts (IB), the removal of uncertainty quantification (UQ), and the full model against the baseline GPT-4o. The results are weighted means across three datasets in Tables 2, Tables 3, and Tables 4. The specific results for each dataset are provided in Table 5 and Table 6 in the Appendix. Removing inductive bias prompts leads to a substantial drop in all generative models. Their consistent results demonstrate that inductive bias is the most important and necessary component when explaining the latent representations of generative models. The removal of uncertainty quantification also results in a slightly decreased performance in all generative models, indicating that uncertainty quantification is also effective, though not as critical as inductive bias prompts. The full model, which incorporates both inductive bias prompts and uncertainty quantification, achieves the highest overall performance, and outperforms all baselines across all models. This confirms the necessity of both inductive bias prompts and uncertainty quantification in our LatentExplainer framework, demonstrating their significant contributions to improving explanatory performance across various generative models. It also shows that our framework can effectively design prompts for different inductive biases in generative models to improve the accuracy of latent explanations.